OpenStudy (anonymous):

Find the remainder when 51^203 is divided by 7

OpenStudy (anonymous):

i don't understand your question

ganeshie8 (ganeshie8):

what have you tried so far ?

OpenStudy (anonymous):

51^203/ 7 = 2^203/7 as, dividing 51 by 7 gives 2 as a remainder! I am stuck there now

ganeshie8 (ganeshie8):

thats a very good start !

OpenStudy (anonymous):

Problem 1.13 is an example

ganeshie8 (ganeshie8):

so you have reduced \(\large 51^{203}/7\) to \(\large 2^{203}/7\) is it ?

OpenStudy (anonymous):

Now what? :-/

ganeshie8 (ganeshie8):

notice that 2^3 / 7 = 8/7 = 1

OpenStudy (anonymous):

and what do we do with 2^200

ganeshie8 (ganeshie8):

\(\large 2^{203} = 2^{3\times 67 + 2} = 2^2*2^{3\times 67} = 4*8^{67}\)

OpenStudy (anonymous):

Well, I did notice that! I didnt get it right

ganeshie8 (ganeshie8):

whats the remainder when u divide 8^67 by 7 ?

OpenStudy (anonymous):

1 ?

ganeshie8 (ganeshie8):

yes, because 8/7 leaves a remainder 1

ganeshie8 (ganeshie8):

\(\large 4*8^{67}/7 = 4*1 = 4\)

OpenStudy (anonymous):

Answer is 4

ganeshie8 (ganeshie8):

yep ! but there is a better notation for remainders. heard of congruence/mod before ?

ganeshie8 (ganeshie8):

you textbook notation is bit misleading - you should switch to below notation which is used by the rest of the world \[\large 2^{203} \equiv 2^{3\times 67 + 2} \equiv 2^2*2^{3\times 67} \equiv 4*8^{67} \equiv 4*1 \equiv 4 \mod 7\]

OpenStudy (anonymous):

Whoaa! Thank you so much :)

ganeshie8 (ganeshie8):

yw :)

OpenStudy (anonymous):

Nope! What is it? Its not the text book! Maybe because I am using my cellphone app to type the queation!